OpenStudy (anonymous):

using the product rule, find the derivative: (x^-5+3)(x^-3+5)

OpenStudy (anonymous):

i keep coming up with 3x^-9 as the first part of the answer but the correct answer is supposed to contain -8x^-9 and I don't see how to arrive at this

OpenStudy (espex):

\[(-5x^{-6})(x^{-3}+5)+(-3^{-4})(x^{-5}+3) \rightarrow -5x^{-9}-25x^{-6}-3x^{-9}-9x^{-4} \] \[\rightarrow -8x^{-9}-25x^{-6}-9x^{-4}\]

OpenStudy (anonymous):

ya so i have all that except for some reason i multiplied exponents and didnt notice thanks a bunch

OpenStudy (espex):

You're welcome.

OpenStudy (anonymous):

can i throw another at you?

OpenStudy (anonymous):

x^2 + 8x + 3 all over / sqrt X I don't get doing square roots with this stuff

OpenStudy (espex):

A square root is just a fractional exponent, an inverse exponent is just a negative exponent. So you can rewrite all of that (if its easier) as \[\frac{x^{2}+8x+3}{\sqrt{x}} \rightarrow \frac{x^{2}+8x+3}{x^{\frac{1}{2}}} \rightarrow (x^{2}+8x+3)(x^{-\frac{1}{2}})\]

OpenStudy (anonymous):

ya so how do i multiply x^2 by x^-1/2, this is the part i dont get

OpenStudy (anonymous):

to me thats just -x

OpenStudy (espex):

When you multiply numbers with exponents you simply add them together, so you would have 2-1/2 or 3/2

OpenStudy (anonymous):

i knew you were going to say that

OpenStudy (espex):

:)

OpenStudy (anonymous):

so then the derivative of x^-1/2 is ????

OpenStudy (espex):

\[x^{-\frac{1}{2}}\rightarrow -\frac{1}{2}x^{-\frac{3}{2}}\rightarrow -\frac{1}{2x^{\frac{3}{2}}}\rightarrow -\frac{1}{2\sqrt{x^{3}}}\]

OpenStudy (anonymous):

thats ridiculous

OpenStudy (anonymous):

so is my answer 3x^2+8x-3 all over 2x^3/2

OpenStudy (espex):

Most likely, I didn't do the math so I will trust that you did. :)

OpenStudy (anonymous):

i obviously suck at math

OpenStudy (anonymous):

thanks for your help, much appreciated

OpenStudy (espex):

Not a problem. :)